# When is an Embedder More Promising than Another?

### v1

@@ Abstract
Embedders play a central role in machine learning, projecting any object into numerical representations that can, in turn, be leveraged to perform various downstream tasks. The evaluation of embedding models typically depends on domain-specific empirical approaches utilizing downstream tasks, primarily because of the lack of a standardized framework for comparison. However, acquiring adequately large and representative datasets for conducting these assessments is not always viable and can prove to be prohibitively expensive and time-consuming. In this paper, we present a unified approach to evaluate embedders. First, we establish theoretical foundations for comparing embedding models, drawing upon the concepts of sufficiency and informativeness. We then leverage these concepts to devise a tractable comparison criterion (information sufficiency), leading to a task-agnostic and self-supervised ranking procedure. We demonstrate experimentally that our approach aligns closely with the capability of embedding models to facilitate various downstream tasks in both natural language processing and molecular biology. This effectively offers practitioners a valuable tool for prioritizing model trials.
@@ 1 Introduction
    A practical relaxation: Estimating deficiency from data samples presents significant challenges. We propose the concept of information sufficiency (IS), which quantifies the information required to simulate one embedder from another (Sec. 3). We estimate the information efficiency to get a task-agnostic and label-free comparison tool for embedders evaluation.
@@ 2.1 Background and notation
We assume that all considered spaces are standard Borel [28] Each such space $\mathsf{U}$ is equipped with its Borel $\sigma$-algebra $\mathcal{B}(\mathsf{U})$. The set of all probability measures on $\mathsf{U}$ is denoted by $\mathcal{P}(\mathsf{U})$ The total variation distance between $P$ and $Q$ is denoted by $\|P-Q\|_{\textrm{TV}}$. Given a joint probability measure $P_{XY}$ induced by two random variables $X\in\mathsf{X}$ and $U\in\mathsf{U}$, the Mutual Information [27] is denoted by $I(X;U)$. A Markov (or transition probability) kernel between $\mathsf{X}$ and $\mathsf{U}$ is a mapping $P_{U|X}:\mathcal{B}(\mathsf{U})\times\mathsf{X}\to[0,1]$. The space of all such $P_{U|X}$ is denoted by $\mathcal{K}(\mathsf{U}|\mathsf{X})$ and $(M\!\circ\!P_{U|X})(Z|x)$ indicates the composition of Markov kernels $M\in\mathcal{K}(\mathsf{Z}|\mathsf{U})$ and $P_{U|X}\in\mathcal{K}(\mathsf{U}|\mathsf{X})$. For further details, refer to Appendix A.
@@ 2.2 Sufficiency and informativeness ordering of embedding models
We aim to compare embedding models without relying on labeled data for downstream tasks. Let us consider two embedding models represented by their Markov kernels (or transition probabilities) $P_{U|X}\in\mathcal{K}(\mathsf{U}|\mathsf{X})$ and $P_{Z|X}\in\mathcal{K}(\mathsf{Z}|\mathsf{X})$, any target set $\mathsf{Y}$ of (discrete or continuous) concepts and feature space $\mathsf{X}$ with joint probability measure $P_{YX}\in\mathcal{P}(\mathsf{Y}\times\mathsf{X})$ induced by random variables $(Y,X)\in\mathsf{Y}\times\mathsf{X}$, as illustrated in Figure 1. First, we study the question:
What sufficient conditions must be met by the embedding model $U$ relative to $Z$ to guarantee that $I(Y;U)\geqslant I(Y;Z)$ for all probability distributions $P_{YX}$
Figure 1: Communicating a concept $y\in\mathsf{Y}$ over two embedding models with prediction $\rho_{Z}(Z)$.
From an information-theoretic perspective [27], the quality of an embedding model can be likened to the capacity of a noisy communication channel with an uncoded input (e.g., a text, a molecule…), where a downstream task of interest is performed at the output (the embedding) of the channel. Let $Y\in\mathsf{Y}$ represent the message (the source) to be communicated over both channels; $X$ represents the transmitted signal; and $P_{U|X}$ and $P_{Z|X}$ the communication channels with outputs $U$ and $Z$, respectively. This process is illustrated in Figure 1. It naturally satisfies the Markov chain $Y\leftrightarrow X\leftrightarrow(U,Z)$. A desirable property is that the embedding models $U$ and $Z$ retain as much pertinent information as feasible to predict $Y$.
We shall be interested in the underlying information relationships between those embedding models that can be interpreted as channel $U$ being "more informative" for communicating $Y$ than channel $Z$. The first attempt to introduce an ordering between communication channels appears in Shannon [94]. Körner and Marton later introduced [57] the concepts of "less noisy" (or more informative) and "degraded" (or sufficiency) orderings between channels.
@@
Let $P_{U|X}$ and $P_{Z|X}$ be two Markov kernels (embedding models).
  • Sufficiency $U\succcurlyeq_{S}Z$. The embedding model $P_{U|X}$ is said to be "sufficient" for the embedding model $P_{U|X}$ (or $Z$ to be degraded w.r.t. $U$) if and only if there exists another Markov kernel $M\in\mathcal{T}(\mathsf{Z}|\mathsf{U})$ such that $\|M\!\circ\!P_{U|X}-P_{Z|X}\|_{\text{TV}}=0$, i.e. $Z$ can be simulated from $U$ using $M$ without information loss).
  • More informative $U\succcurlyeq_{I}Z$. The embedding model $P_{U|X}$ is said to be "more informative" (or less noisy) than $P_{Z|X}$ if and only if the embedding models always satisfy the inequality
    $I(Y;U)\geqslant I(Y;Z),\quad\forall P_{YX}\in\mathcal{P}(\mathsf{Y}\times\mathsf{X}).$
@@
    Informativeness $\Rightarrow$ higher capacity to distinguish concepts. If the embedding model $P_{U|X}$ is more informative than embedding model $P_{Z|X}$, i.e. $U\succcurlyeq_{I}Z$, then
    $\textrm{KL}\big{(}P_{U|Y}(\cdot|y_{0})\|P_{U|Y}(\cdot|y_{1})\big{)}\geqslant\textrm{KL}\big{(}P_{Z|Y}(\cdot|y_{0})\|P_{Z|Y}(\cdot|y_{1})\big{)},$
@@
An immediate consequence of claim (i) is that the sufficient condition between embedding models implies that the embedding model $U$ is more informative than the embedding model $Z$ relative to all target concepts in $\mathsf{Y}$ over all possible data distributions: $I(Y;U)\geqslant I(Y;Z)$, for all probability distributions $P_{YX}$.
Although $U$ being more informative than $Z$ does not necessarily imply $U\succcurlyeq_{S}Z$ [57, 66]; (ii) states that being more informative ensures a higher statistical discrimination capacity between any pairs of target concepts (for further discussion, see Sec. B.2).
Motivated by the concepts of sufficiency and informativeness between embedding models, we can inquire about their statistical consequences for a learner conducting an inference task on these embeddings. More precisely, given a finite set of concepts $\mathsf{Y}$, if $U\succcurlyeq_{S}Z$, is the Bayes risk expected to be smaller when the inference is based on $U$ than when it is based on $Z$?
@@ 2.3 Comparing statistical experiments with embedding Models
In our framework, a statistical experiment [13] consists of a mathematical abstraction (see Appendix A for further details) intended to represent a downstream task where a learner aims at inferring a concept $y\in\mathsf{Y}$ from the embeddings $U$ or $Z$. Deciding what embedder should be used to perform a given task is too general. In this work, we do not take into account the computational cost or the size of an embedder and solely focus on the following question:
What are the necessary and sufficient conditions that ensure that employing the embedding $U$ for any task $P_{YX}$ leads to lower risk compared to using the embedding $Z$?
@@
Given two embedding models $P_{U|X}\in\mathcal{K}(\mathsf{U}|\mathsf{X})$ and $P_{Z|X}\in\mathcal{K}(\mathsf{Z}|\mathsf{X})$, the following statements are equivalent:
@@
    The embedding model $P_{U|X}$ is sufficient relative to $P_{Z|X}$, i.e. $U\succcurlyeq_{S}Z$.
@@
    For all probability measures $P_{YX}$ on finite alphabets $\mathsf{Y}\times\mathsf{X}$, the expected Bayes risks satisfy
    $\inf_{\rho_{U}:\mathsf{U}\rightarrow\mathcal{P}(\mathsf{Y})}\Pr\big{(}\rho_{Y}(U)\neq Y\big{)}\leqslant\inf_{\rho_{Z}:\mathsf{Z}\rightarrow\mathcal{P}(\mathsf{Y})}\Pr\big{(}\rho_{Z}(Z)\neq Y\big{)}.$
@@
In other words, if we can fully simulate an embedder $Z$ from another embedder $U$, the expected risk across all potential classification tasks cannot be greater when using $U$ compared to $Z$. The proof of this Proposition is given in Sec. B.3. It is worth mentioning that various versions of this result are available in the literature [104]. However, our extension here, in a simpler setting, incorporates concepts and features into the experiment comparison framework.
@@ 2.4 Challenges in ranking embedding models and their deficiency
  • Equivalence: $U\succcurlyeq_{S}Z$ and $Z\succcurlyeq_{S}U$ denoted $U\approx Z$. U and Z can simulate each other.
  • Comparability: $U\succcurlyeq_{S}Z$ but $Z\not\succcurlyeq_{S}U$ only Z can be simulated from U.
  • Non-comparability: $U\not\succcurlyeq_{S}Z$ and $Z\not\succcurlyeq_{S}U$, neither U nor Z can simulate each other.
@@
The deficiency $\delta(P_{U|X}\rightarrow P_{Z|X})$ of $P_{Z|X}$ relative to $P_{U|X}$ is defined as [63]
$\delta(P_{U|X}\rightarrow P_{Z|X})\triangleq\inf_{M\in\mathcal{K}(\mathsf{Z}|\mathsf{U})}\|M\!\circ\!P_{U|X}-P_{Z|X}\|_{\text{TV}},$
where the infimum is taken over all Markov kernels (or transition probabilities) $M\in\mathcal{K}(\mathsf{Z}|\mathsf{U})$, mapping stochastically $\mathsf{U}$ and $\mathsf{Z}$, and $\delta$ measures error between the simulated and true embedders.
$\delta$ indicates how well one model can be reconstructed from the other, it induces a natural relaxation of the sufficiency where the reconstruction does not have to be perfect [^1] If $U\succcurlyeq_{S},Z$, then $\delta(P_{Z|X}\rightarrow P_{U|X})=0$, while if $U\not\succcurlyeq_{S}Z$, then $\delta(P_{Z|X}\rightarrow P_{U|X})>0$. for us to obtain guarantees on the downstream tasks performance (See Corollary 1). It avoids the non-comparability problem by evaluating "how much information" we lose when passing from one model to the other one.
Le Cam [63] showed that, for a given task $Y$, the deficiency $\delta(P_{U|Y}\rightarrow P_{Z|Y})$ is directly related to the expected Bayes risks on the task (see Sec. B.6). We extend this result to the comparison of two embedding models $P_{U|X}$ and $P_{Z|X}$ in a task-agnostic manner and build the relation to the expected Bayes risks for any classification task $Y$, defined as $\mathcal{R}_{U}(Y)\triangleq\inf\limits_{\rho_{U}:\mathsf{U}\rightarrow\mathcal{P}(\mathsf{Y})}\Pr\big{(}\rho_{U}(U)\neq Y\big{)}$.
@@
Given two embedding models $P_{U|X}$ and $P_{Z|X}$, the following statements are equivalent:
@@
    The deficiency $\delta(P_{U|X}\rightarrow P_{Z|X})\leqslant\epsilon$.
@@
    For any probability distribution $P_{YX}$ on finite alphabets $\mathsf{Y}\times\mathsf{X}$, there exists $\epsilon>0$ satisfying
    $\mathcal{R}_{U}(Y)-\epsilon|\mathsf{X}|\leqslant\mathcal{R}_{Z}(Y).$
@@
In particular, we can infer that for any classification task $Y$, the expected Bayes risk of the embedding model $U$ is upper bounded by the expected Bayes risk of the embedding model $Z$:
$\mathcal{R}_{U}(Y)-\mathcal{R}_{Z}(Y)\leqslant|\mathsf{X}|\delta(P_{U|X}\rightarrow P_{Z|X}),\quad\textrm{for all distributions $P_{YX}$, }$
and similarly, $|\mathcal{R}_{U}(Y)-\mathcal{R}_{Z}(Y)|\leqslant|\mathsf{X}|\max\big{\{}\delta(P_{U|X}\rightarrow P_{Z|X}),\delta(P_{Z|X}\rightarrow P_{U|X})\big{\}}$, for all distributions $P_{YX}$. If both deficiencies are small, the resulting expected Bayes risks of the embedding models $U$ and $Z$ will be close to each other for any target task $Y$.

### v2

@@ Abstract
Embedders play a central role in machine learning, projecting any object into numerical representations that can, in turn, be leveraged to perform various downstream tasks. The evaluation of embedding models typically depends on domain-specific empirical approaches utilizing downstream tasks, primarily because of the lack of a standardized framework for comparison. However, acquiring adequately large and representative datasets for conducting these assessments is not always viable and can prove to be prohibitively expensive and time-consuming. In this paper, we present a unified approach to evaluate embedders. First, we establish theoretical foundations for comparing embedding models, drawing upon the concepts of sufficiency and informativeness. We then leverage these concepts to devise a tractable comparison criterion (information sufficiency), leading to a task-agnostic and self-supervised ranking procedure. We demonstrate experimentally that our approach aligns closely with the capability of embedding models to facilitate various downstream tasks in both natural language processing and molecular biology. This effectively offers practitioners a valuable tool for prioritizing model trials. [^1] The code used to perform all experiments is available at https://github.com/ills-montreal/emir
@@ 1 Introduction
    A practical relaxation: Estimating deficiency presents significant challenges. We propose the concept of information sufficiency (IS), which quantifies the information required to simulate one embedder from another (Sec. 3). We estimate the information efficiency to get a task-agnostic and label-free comparison tool for embedders evaluation.
@@ 2.1 Background and notation
We assume that all considered spaces are standard Borel [28] Each such space $\mathsf{U}$ is equipped with its Borel $\sigma$-algebra $\mathcal{B}(\mathsf{U})$. The set of all probability measures on $\mathsf{U}$ is denoted by $\mathcal{P}(\mathsf{U})$ The total variation distance between $P$ and $Q$ is denoted by $\|P-Q\|_{\textrm{TV}}$. Given a joint probability measure $P_{XY}$ induced by two random variables $X\in\mathsf{X}$ and $U\in\mathsf{U}$, the Mutual Information [27] is denoted by $I(X;U)$. A Markov (or transition probability) kernel between $\mathsf{X}$ and $\mathsf{U}$ is a mapping $P_{U|X}:\mathcal{B}(\mathsf{U})\times\mathsf{X}\to[0,1]$. The space of all such $P_{U|X}$ is denoted by $\mathcal{K}(\mathsf{U}|\mathsf{X})$ and $(M\!\circ\!P_{U|X})(V|x)$ indicates the composition of Markov kernels $M\in\mathcal{K}(\mathsf{V}|\mathsf{U})$ and $P_{U|X}\in\mathcal{K}(\mathsf{U}|\mathsf{X})$. For further details, refer to Appendix A.
@@ 2.2 Sufficiency and informativeness ordering of embedding models
We aim to compare embedding models without relying on labeled data for downstream tasks. Let us consider two embedding models represented by their Markov kernels (or transition probabilities) $P_{U|X}\in\mathcal{K}(\mathsf{U}|\mathsf{X})$ and $P_{V|X}\in\mathcal{K}(\mathsf{V}|\mathsf{X})$, any target set $\mathsf{Y}$ of (discrete or continuous) concepts and feature space $\mathsf{X}$ with joint probability measure $P_{YX}\in\mathcal{P}(\mathsf{Y}\times\mathsf{X})$ induced by random variables $(Y,X)\in\mathsf{Y}\times\mathsf{X}$, as illustrated in Figure 1. First, we study the question:
What sufficient conditions must be met by the embedding model $U$ relative to $V$ to guarantee that $I(Y;U)\geqslant I(Y;V)$ for all distributions $P_{YX}$
Figure 1: Communicating a concept $y\in\mathsf{Y}$ over two embedding models with prediction $\rho_{V}(V)$.
From an information-theoretic perspective [27], the quality of an embedding model can be likened to the capacity of a noisy communication channel with an uncoded input (e.g., a text, a molecule…), where a downstream task of interest is performed at the output (the embedding) of the channel. Let $Y\in\mathsf{Y}$ represent the message (the source) to be communicated over both channels; $X$ represents the transmitted signal; and $P_{U|X}$ and $P_{V|X}$ the communication channels with outputs $U$ and $V$, respectively. This process is illustrated in Figure 1. It naturally satisfies the Markov chain $Y\leftrightarrow X\leftrightarrow(U,V)$. A desirable property is that the embedding models $U$ and $V$ retain as much pertinent information as feasible to predict $Y$.
We shall be interested in the underlying information relationships between those embedding models that can be interpreted as channel $U$ being "more informative" for communicating $Y$ than channel $V$. The first attempt to introduce an ordering between communication channels appears in Shannon [94]. Körner and Marton later introduced [57] the concepts of "less noisy" (or more informative) and "degraded" (or sufficiency) orderings between channels.
@@
Let $P_{U|X}$ and $P_{V|X}$ be two Markov kernels (embedding models).
  • Sufficiency $U\succcurlyeq_{S}V$. The embedding model $P_{U|X}$ is said to be "sufficient" for the embedding model $P_{U|X}$ (or $V$ to be degraded w.r.t. $U$) if and only if there exists another Markov kernel $M\in\mathcal{T}(\mathsf{V}|\mathsf{U})$ such that $\mathbb{E}\|M\!\circ\!P_{U|X}-P_{V|X}\|_{\text{TV}}=0$, i.e. $V$ can be simulated from $U$ using $M$ without information loss).
  • More informative $U\succcurlyeq_{I}V$. The embedding model $P_{U|X}$ is said to be "more informative" (or less noisy) than $P_{V|X}$ if and only if the embedding models always satisfy the inequality
    $I(Y;U)\geqslant I(Y;V),\quad\forall P_{YX}\in\mathcal{P}(\mathsf{Y}\times\mathsf{X}).$
@@
    Informativeness $\Rightarrow$ higher capacity to distinguish concepts. If the embedding model $P_{U|X}$ is more informative than embedding model $P_{V|X}$, i.e. $U\succcurlyeq_{I}V$, then
    $\textrm{KL}\big{(}P_{U|Y}(\cdot|y_{0})\|P_{U|Y}(\cdot|y_{1})\big{)}\geqslant\textrm{KL}\big{(}P_{V|Y}(\cdot|y_{0})\|P_{V|Y}(\cdot|y_{1})\big{)},$
@@
An immediate consequence of claim (i) is that the sufficient condition between embedding models implies that the embedding model $U$ is more informative than the embedding model $V$ relative to all target concepts in $\mathsf{Y}$ over all possible data distributions: $I(Y;U)\geqslant I(Y;V)$, for all probability distributions $P_{YX}$.
Although $U$ being more informative than $V$ does not necessarily imply $U\succcurlyeq_{S}V$ [57, 66]; (ii) states that being more informative ensures a higher statistical discrimination capacity between any pairs of target concepts (for further discussion, see Sec. B.2).
Motivated by the concepts of sufficiency and informativeness between embedding models, we can inquire about their statistical consequences for a learner conducting an inference task on these embeddings. More precisely, given a finite set of concepts $\mathsf{Y}$, if $U\succcurlyeq_{S}V$, is the Bayes risk expected to be smaller when the inference is based on $U$ than when it is based on $V$?
@@ 2.3 Comparing statistical experiments with embedding Models
In our framework, a statistical experiment [13] consists of a mathematical abstraction (see Appendix A for further details) intended to represent a downstream task where a learner aims at inferring a concept $y\in\mathsf{Y}$ from the embeddings $U$ or $V$. Deciding what embedder should be used to perform a given task is too general. In this work, we do not take into account the computational cost or the size of an embedder and solely focus on the following question:
What are the necessary and sufficient conditions that ensure that employing the embedding $U$ for any task $P_{YX}$ leads to lower risk compared to using the embedding $V$?
@@
Given two embedding models $P_{U|X}\in\mathcal{K}(\mathsf{U}|\mathsf{X})$ and $P_{V|X}\in\mathcal{K}(\mathsf{V}|\mathsf{X})$, the following statements are equivalent:
@@
    The embedding model $P_{U|X}$ is sufficient relative to $P_{V|X}$, i.e. $U\succcurlyeq_{S}V$.
@@
    For all conditional probability measures $P_{Y|X}$ on finite alphabet $\mathsf{Y}$, the Bayes risks satisfy
    $\inf_{\rho_{U}:\mathsf{U}\rightarrow\mathcal{P}(\mathsf{Y})}\Pr\big{(}\hat{Y}_{U}\neq Y\big{)}\leqslant\inf_{\rho_{V}:\mathsf{V}\rightarrow\mathcal{P}(\mathsf{Y})}\Pr\big{(}\hat{Y}_{V}\neq Y\big{)},$
    where $\hat{Y}_{U}$ and $\hat{Y}_{V}$ are distributed according to $\rho_{U}(U)$ and $\rho_{V}(V)$, respectively.
@@
In other words, if we can fully simulate an embedder $V$ from another embedder $U$, the expected risk across all potential classification tasks cannot be greater when using $U$ compared to $V$. The proof of this Proposition is given in Sec. B.3. It is worth mentioning that various versions of this result are available in the literature [104]. However, our extension here, in a simpler setting, incorporates concepts and features into the experiment comparison framework.
@@ 2.4 Challenges in ranking embedding models and their deficiency
  • Equivalence: $U\succcurlyeq_{S}V$ and $V\succcurlyeq_{S}U$ denoted $U\approx V$; $U$ and $V$ can simulate each other.
  • Comparability: $U\succcurlyeq_{S}V$ but $V\not\succcurlyeq_{S}U$ only $V$ can be simulated from $U$.
  • Non-comparability: $U\not\succcurlyeq_{S}V$ and $V\not\succcurlyeq_{S}U$, neither $U$ nor $V$ can simulate each other.
@@
The deficiency $\delta(P_{U|X}\rightarrow P_{V|X})$ of $P_{V|X}$ relative to $P_{U|X}$ is defined as [63]
$\delta(P_{U|X}\rightarrow P_{V|X})\triangleq\inf_{M\in\mathcal{K}(\mathsf{V}|\mathsf{U})}\mathbb{E}\|M\!\circ\!P_{U|X}-P_{V|X}\|_{\text{TV}},$
where the infimum is taken over all Markov kernels (or transition probabilities) $M\in\mathcal{K}(\mathsf{V}|\mathsf{U})$, mapping stochastically $\mathsf{U}$ and $\mathsf{V}$, and $\delta$ measures error between the simulated and true embedders.
$\delta$ indicates how well one model can be reconstructed from the other, it induces a natural relaxation of the sufficiency where the reconstruction does not have to be perfect [^2] If $U\succcurlyeq_{S},V$, then $\delta(P_{V|X}\rightarrow P_{U|X})=0$, while if $U\not\succcurlyeq_{S}V$, then $\delta(P_{V|X}\rightarrow P_{U|X})>0$. for us to obtain guarantees on the downstream tasks performance (See Corollary 1). It avoids the non-comparability problem by evaluating "how much information" we lose when passing from one model to the other one.
Le Cam [63] showed that, for a given task $Y$, the deficiency $\delta(P_{U|Y}\rightarrow P_{V|Y})$ is directly related to the expected Bayes risks on the task (see Sec. B.6). We extend this result to the comparison of two embedding models $P_{U|X}$ and $P_{V|X}$ in a task-agnostic manner and build the relation to the expected Bayes risks for any classification task $Y$.
@@
Given two embedding models $P_{U|X}$ and $P_{V|X}$ satisfying:
@@
    The deficiency $\delta(P_{U|X}\rightarrow P_{V|X})\leqslant\gamma$.
@@
    For any conditional distribution $P_{Y|X}$ on finite alphabets $\mathsf{Y}$,
    $\inf_{\rho_{U}:\mathsf{U}\rightarrow\mathcal{P}(\mathsf{Y})}\Pr\big{(}\hat{Y}_{U}\neq Y\big{)}-\varepsilon\leqslant\inf_{\rho_{V}:\mathsf{V}\rightarrow\mathcal{P}(\mathsf{Y})}\Pr\big{(}\hat{Y}_{V}\neq Y\big{)}.$
Statement (ii) implies (i) provided that $\gamma\geqslant 2|\mathsf{Y}|\varepsilon$ and conversely, (i) implies (ii) provided that $\gamma\leqslant\varepsilon$.
@@
In particular, we can infer that for any classification task $Y$, the expected Bayes risk of the embedding model $U$, denoted by $\mathcal{R}_{U}$, is upper bounded by the expected Bayes risk of the embedding model $V$, denoted by $\mathcal{R}_{V}$:
$\mathcal{R}_{U}-\mathcal{R}_{V}\leqslant\delta(P_{U|X}\rightarrow P_{V|X}),\quad\textrm{for all conditional distributions $P_{Y|X}$, }$
and similarly, $|\mathcal{R}_{U}-\mathcal{R}_{V}|\leqslant\max\big{\{}\delta(P_{U|X}\rightarrow P_{V|X}),\delta(P_{V|X}\rightarrow P_{U|X})\big{\}}$, for all conditional distributions $P_{Y|X}$. If both deficiencies are small, the resulting expected Bayes risks of the embedding models $U$ and $V$ will be close to each other for any target task $Y$.